You are managing the investments of a pension fund. You have to make a payment to your customers in 6 months. The total amount of the payments is equal to $2,000,000 . How much do you have to invest each month, starting next month, for 5 months to exactly provide for the payments if your investments earn 5.00% APR (compounded monthly)?

Calculate your answer to the nearest dollar. Write only the number with no dollar sign or comma (e.g., 3711)
Answer #1

Rate per month =5%/12
Number of years =5
FV =2000000
Monthly Payment =FV/((1+r)^n-1)/r) =2000000/(((1+5%/12)^5-1)/(5%/12))) =396680.53 or 396681